# Do Test Scores Help Teachers Give Better Track Advice to Students?

### v1

@@ 5 An alternative to assuming the Exclusion Restriction
Figure 3 compares, for the three aggregate tracks and for the H, AH and AL strata, the point identified $APCE_{J}$ estimates obtained under unconfoundedness (Assumption 6, red dot), that are derived in the Online Appendix to Section 3.4; the bounds of the $APCE_{J}$ obtained under the ER (Assumption 3, solid black), that are derived in Section 3; and the bounds of the $APCE_{J}$ obtained under PI and Homogeneity (Assumptions 8 and 7, dashed light-blue), that are derived in this section. In the case of H and AL students, unconfoundedness delivers estimates that are located near the lower bounds of the partially identified alternative estimates. For the AH students, both with and without unconfoundedness, the $APCE_{J}$ estimates are precise and close one to the other.
@@ 6 Fairness of teachers’ recommendations
According to this definition, a recommender is fair as long as her recommendations are independent of the protected attribute among students in the same stratum. The analogous fraction in other strata can be lower or higher if it is equal for both genders within each stratum. A test for Principal Fairness of a source of secondary school track recommendations can then be designed as follows. Let the two sources of recommendations that we would like to compare be: “teacher alone” ($z=0$) versus “teachers informed by test scores” ($z=1$). Given two values $b$ and $b^{\prime}$ of a protected attribute $B_{i}$, the Principal Fairness of source $z$ in stratum $j$ is given by:
@@
The sign of $\Delta_{j}(z)$ is also important, as it indicates the direction of discrimination. Suppose again that $B_{i}=b$ denotes immigrant students. Then finding, for example, that $\Delta_{j}(1)>0$ while $\Delta_{j}(0)<0$, would mean that when test score information is provided to recommenders, immigrant students are positively discriminated into the high track, while when the recommenders decide without this information, immigrants are negatively discriminated. Note that in our setting, where teachers are officially allowed to upgrade only students who score above the cutoff, all the $\Delta_{j}(0)$ should be equal to 0 by definition. Therefore, a $\Delta_{j}(0)=0$ cannot be interpreted as evidence of fairness in the provisional track recommendation made by teachers without the test scores. On the other hand, a $\Delta_{j}(0)$ different from 0 would indicate different rates of non-compliance below the cutoff across protected attributes in stratum $j$.
@@ 7 Conclusions
Second, our results show that information based on standardized test scores can assist teachers in giving better recommendations, and its provision should, therefore, be considered.

### v2

@@ 5 An alternative to assuming the Exclusion Restriction
Figure 3 compares, for the three aggregate tracks and for the H, AH, and AL strata, the point identified $APCE_{J}$ estimates obtained under unconfoundedness (Assumption 6, red dot), that are derived in the Online Appendix to Section 3.4; the bounds of the $APCE_{J}$ obtained under the ER (Assumption 3, solid black), that are derived in Section 3; and the bounds of the $APCE_{J}$ obtained under PI and Homogeneity (Assumptions 8 and 7, dashed light-blue), that are derived in this section. In the case of H and AL students, unconfoundedness delivers estimates that are located near the lower bounds of the partially identified alternative estimates. For the AH students, both with and without unconfoundedness, the $APCE_{J}$ estimates are precise and close one to the other.
@@ 6 Fairness of teachers’ recommendations
According to this definition, a recommender is fair as long as her recommendations are independent of the protected attribute among students in the same stratum. The analogous fraction in other strata can be lower or higher if it is equal for both genders within each stratum. A test for Principal Fairness of a source of secondary school track recommendations can then be designed as follows. Let the two sources of recommendations that we would like to compare be: “teacher alone” ($z=0$) versus “teacher informed by test scores and allowed to upgrade” ($z=1$). Given two values $b$ and $b^{\prime}$ of a protected attribute $B_{i}$, the Principal Fairness of source $z$ in stratum $j$ is given by:
@@
The sign of $\Delta_{j}(z)$ is also important, as it indicates the direction of discrimination. Suppose again that $B_{i}=b$ denotes immigrant students. Then finding, for example, that $\Delta_{j}(1)>0$ while $\Delta_{j}(0)<0$, would mean that when recommenders are allowed to upgrade their advice based on test scores, immigrant students are positively discriminated into the high track, while when the recommenders give advice without seeing test scores, immigrants are negatively discriminated. Note that in our setting, where teachers are officially allowed to upgrade only students who score above the cutoff, all the $\Delta_{j}(0)$ should be equal to 0 by definition. Therefore, a $\Delta_{j}(0)=0$ cannot be interpreted as evidence of fairness in the provisional track recommendation made by teachers without the test scores. On the other hand, a $\Delta_{j}(0)$ different from 0 would indicate different rates of non-compliance below the cutoff across protected attributes in stratum $j$.
@@ 7 Conclusions
Second, our results show that allowing teachers to upgrade their recommendations based on standardized test scores helps them to improve the quality of track advising. Ideally, one would like to compare recommendations given by teachers with and without the information provided by test scores, which is not what the quasi-experiment we study can offer. However, our evidence suggests that test scores do help teachers provide better recommendations in general, not just to upgrade previous ones. Therefore, appropriate controlled experiments should be implemented to further study and formally test the more general hypothesis.